OpenStudy (anonymous):

5) A 40.0-kg wagon is towed up a hill inclined at 18.5º with respect to the horizontal. The tow rope is parallel to the incline and has a tension of 140 N in it. Assume that the wagon starts from rest at the bottom of the hill, and neglect friction. How fast is the wagon moving after moving 80.0 m up the hill?

OpenStudy (anonymous):

have you learned about rotating the x and y axis?

OpenStudy (anonymous):



OpenStudy (anonymous):

if we rotate the the incline down, so it is flat, it will make our math much easier, so like this 

OpenStudy (anonymous):

so, now we sum the forces in the x direction +140 - cos(18.5)*40*9.81 = 5.7 since it's 5.7, we know there is accelearation! and remember F = m*a, then 5.7 = ma now what is accelerating? the wagon of 40kg so 5.7 = 40*a, solve for a

OpenStudy (anonymous):

@TheLacey you still here?